# Distribution Matching for Multi-Task Learning of Classification Tasks: a Large-Scale Study on Faces & Beyond

### 02 Jan 2024 v1

@@ Introduction
Up until now training holistic models has been primarily addressed by combining multiple datasets to solve individual tasks (Ranjan et al. 2017), or by collecting the annotations in terms of all tasks (Zamir et al. 2018; Kokkinos 2017; Arsenos, Kollias, and Kollias 2022; Kollias, Arsenos, and Kollias 2023c; Kollias et al. 2018b). For example, in affective computing, two most common tasks are predicting categorical expressions (e.g., happy, sad) and activations of binary action units (Ekman 1997) (activation of facial muscles) to explain the affective state. Collecting annotations of AUs is particularly costly, as it requires skilled annotators. The datasets collected so far (Mollahosseini, Hasani, and Mahoor 2017; Benitez-Quiroz, Srinivasan, and Martinez 2016) have annotations for training only one task and, despite significant effort, there is no dataset that for has complete annotations of both tasks. Co-training via task relatedness is an effective way of aggregating knowledge across datasets and transferring it across tasks, especially with little or non-overlapping annotations, or when not many training data are available, or when there is a big discrepancy in the size of labeled data per task.
@@
Works exist in literature that use expression labels to complement missing AU annotations or increase generalization of AU classifiers (Yang et al. 2016; Wang, Gan, and Ji 2017; Kollias et al. 2023b, a). Our work deviates from such methods, as we target joint learning of both facial behavior tasks via a single framework, whilst these works perform only AU detection. In the face analysis domain, the use of MTL is somewhat limited. In (Wang et al. 2017), MTL was tackled through a network that jointly handled face recognition and facial attribute prediction tasks. At first a network was trained for facial attribute detection; then it was used to generate attribute labels for another database that only contained face identification labels. Then, a MT network was trained using that database. The network’s loss function was the sum of the independent task losses. In (Deng, Chen, and Shi 2020), a unified model performing facial AU detection, expression recognition, and valence-arousal estimation was proposed; the utilized database contained images not annotated for all tasks. To tackle this, authors trained a teacher MT model using the complete labels, then generated pseudo-annotations and finally trained a student MT model on the union of original and pseudo labels; that network outperformed the teacher one. The teacher model did not exploit the tasks being interconnected - the overall loss was equal to the sum of the independent task losses. Thus, the student model did not learn this relatedness. This work utilized only one database.
@@
2) Inferred Empirically from Dataset Annotations: If the above cognitive study is not available, we can infer task relatedness from external dataset annotations. In particular, we use the training set of Aff-Wild2 database (Kollias and Zafeiriou 2019, 2021a, 2021b; Kollias et al. 2017a, 2019, 2020b, 2023b; Zafeiriou et al. 2017; Kollias, Sharmanska, and Zafeiriou 2019, 2021; Kollias and Zafeiriou 2020) to infer task relatedness, since this dataset is the first in-the-wild one that is fully annotated with basic expressions and AUs; this is shown in Table 1. In the following, we use as domain knowledge the cognitive and psychological study (Du, Tao, and Martinez 2014), to encode task relatedness and introduce the proposed approach for coupling the tasks.
@@
Here, we show that our approach can also be used in other application scenarios: i) face recognition (facial attribute detection and face identification); ii) fine-grained species categorization (species classification and attribute detection); iii) shoe type recognition (shoe type classification and attribute detection); iv) clothing categories recognition (classification of clothing categories and attributes) (Kollias et al. 2021; Bouas et al. 2020; Kollias et al. 2020c; Kollias and Zafeiriou 2018a).
@@ C MT (Aff-Wild2)
In this work we utilized: AffectNet (Mollahosseini, Hasani, and Mahoor 2017) with around 350K in-the-wild images annotated for 7 basic expressions; RAF-DB (Li, Deng, and Du 2017) with around 15K in-the-wild images annotated for 7 basic expressions; ABAW4 LSD (Kollias 2022a, 2023; Kollias et al. 2023c) -utilized in 4th Affective Behavior Analysis in-the-wild (ABAW) Competition at ECCV 2022- with around 280K in-the-wild synthetic images and 100K in-the-wild real images (which constitute the test set) annotated for 6 basic expressions; Aff-Wild2 (Kollias 2022b; Kollias, Arsenos, and Kollias 2023a; Kollias et al. 2018a; Kollias and Zafeiriou 2018b) -as utilized in 3rd ABAW Competition at CVPR 2022- with 564 in-the-wild videos (A/V) of around 2.8M frames annotated for 7 basic expressions (plus ’other’), 12 AUs and valence-arousal; EmotioNet (Fabian Benitez-Quiroz, Srinivasan, and Martinez 2016) with around 50K images manually annotated for 11 AUs; CelebA with around 205K in-the-wild images of around 10.2K identities, each with 40 attributes (its training, validation and test sets are subject independent; for our experiments, we generated a new split into 3 subject dependent sets); Caltech-UCSD Birds-200-2011 (Wah et al. 2011) (CUB) with around 12K images of 200 bird species and of 312 binary attributes; Shoes (Wah et al. 2011) (S-ADD) with around 15K women shoe images of 10 different types and of 10 attributes (Kovashka, Parikh, and Grauman 2012) (it does not contain any predefined split, thus we split it in a training set of 7.4K and a test set of 7.3K images). Clothing Attributes Dataset (Chen, Gallagher, and Girod 2012) (CAD) with around 2K images partially annotated for 7 clothing categories, 23 binary and 2 multi-class attributes (due to its very small size and to the non-predefined split, we perform 6 times 2-fold cross validation, i.e., we create 6 different 50-50 splits of the data).
@@
Case Study I: We used RetinaFace (Deng et al. 2020) to extract bboxes and 5 facial landmarks (for alignment) and resized images to $112\times 112\times 3$. Mixaugment (Psaroudakis and Kollias 2022) was used for data augmentation. Case Study II: for CelebA we used the database’s aligned images and resized them to $112\times 112\times 3$ (batch size = 200, Adam, lr = $10^{-3}$); for CUB, we cropped the images using the bboxes, resized them to $280\times 280\times 3$, we used label smoothing (value = 0.3) and performed affine transformations as data augmentation (batch size = 150, Adam, lr = $10^{-3}$); for S-ADD, we resized images to $280\times 280\times 3$ (batch size = 100, Adam, lr = $10^{-3}$); for CAD, we resized images to $280\times 280\times 3$ (batch size = 50, Adam, lr = $10^{-3}$). In all databases, images were normalized to $[-1,1]$ (Kollias, Arsenos, and Kollias 2023b; Kollias, Tagaris, and Stafylopatis 2016; Kollias et al. 2017b, 2015; Salpea, Tzouveli, and Kollias 2023; Papaioannou et al. 2022; Arsenos et al. 2023). Tesla V100 32GB GPU & Tensorflow were used for model training.

### 03 Jan 2024 v2

@@ Introduction
Up until now training holistic models has been primarily addressed by combining multiple datasets to solve individual tasks (Ranjan et al. 2017), or by collecting the annotations in terms of all tasks (Zamir et al. 2018; Kokkinos 2017). For example, in affective computing, two most common tasks are predicting categorical expressions (e.g., happy, sad) and activations of binary action units (Ekman 1997) (activation of facial muscles) to explain the affective state. Collecting annotations of AUs is particularly costly, as it requires skilled annotators. The datasets collected so far (Mollahosseini, Hasani, and Mahoor 2017; Benitez-Quiroz, Srinivasan, and Martinez 2016) have annotations for training only one task and, despite significant effort, there is no dataset that for has complete annotations of both tasks. Co-training via task relatedness is an effective way of aggregating knowledge across datasets and transferring it across tasks, especially with little or non-overlapping annotations, or when not many training data are available, or when there is a big discrepancy in the size of labeled data per task.
@@
Works exist in literature that use expression labels to complement missing AU annotations or increase generalization of AU classifiers (Yang et al. 2016; Wang, Gan, and Ji 2017). Our work deviates from such methods, as we target joint learning of both facial behavior tasks via a single framework, whilst these works perform only AU detection. In the face analysis domain, the use of MTL is somewhat limited. In (Wang et al. 2017), MTL was tackled through a network that jointly handled face recognition and facial attribute prediction tasks. At first a network was trained for facial attribute detection; then it was used to generate attribute labels for another database that only contained face identification labels. Then, a MT network was trained using that database. The network’s loss function was the sum of the independent task losses. In (Deng, Chen, and Shi 2020), a unified model performing facial AU detection, expression recognition, and valence-arousal estimation was proposed; the utilized database contained images not annotated for all tasks. To tackle this, authors trained a teacher MT model using the complete labels, then generated pseudo-annotations and finally trained a student MT model on the union of original and pseudo labels; that network outperformed the teacher one. The teacher model did not exploit the tasks being interconnected - the overall loss was equal to the sum of the independent task losses. Thus, the student model did not learn this relatedness. This work utilized only one database.
@@
2) Inferred Empirically from Dataset Annotations: If the above cognitive study is not available, we can infer task relatedness from external dataset annotations. In particular, we use the training set of Aff-Wild2 database (Kollias and Zafeiriou 2019, 2021a, 2021b; Kollias et al. 2017, 2019, 2020b, 2023; Zafeiriou et al. 2017) to infer task relatedness, since this dataset is the first in-the-wild one that is fully annotated with basic expressions and AUs; this is shown in Table 1. In the following, we use as domain knowledge the cognitive and psychological study (Du, Tao, and Martinez 2014), to encode task relatedness and introduce the proposed approach for coupling the tasks.
@@
Here, we show that our approach can also be used in other application scenarios: i) face recognition (facial attribute detection and face identification); ii) fine-grained species categorization (species classification and attribute detection); iii) shoe type recognition (shoe type classification and attribute detection); iv) clothing categories recognition (classification of clothing categories and attributes).
@@ C MT (Aff-Wild2)
In this work we utilized: AffectNet (Mollahosseini, Hasani, and Mahoor 2017) with around 350K in-the-wild images annotated for 7 basic expressions; RAF-DB (Li, Deng, and Du 2017) with around 15K in-the-wild images annotated for 7 basic expressions; ABAW4 LSD (Kollias 2022a) -utilized in 4th Affective Behavior Analysis in-the-wild (ABAW) Competition at ECCV 2022- with around 280K in-the-wild synthetic images and 100K in-the-wild real images (which constitute the test set) annotated for 6 basic expressions; Aff-Wild2 (Kollias 2022b) -as utilized in 3rd ABAW Competition at CVPR 2022- with 564 in-the-wild videos (A/V) of around 2.8M frames annotated for 7 basic expressions (plus ’other’), 12 AUs and valence-arousal; EmotioNet (Fabian Benitez-Quiroz, Srinivasan, and Martinez 2016) with around 50K images manually annotated for 11 AUs; CelebA with around 205K in-the-wild images of around 10.2K identities, each with 40 attributes (its training, validation and test sets are subject independent; for our experiments, we generated a new split into 3 subject dependent sets); Caltech-UCSD Birds-200-2011 (Wah et al. 2011) (CUB) with around 12K images of 200 bird species and of 312 binary attributes; Shoes (Wah et al. 2011) (S-ADD) with around 15K women shoe images of 10 different types and of 10 attributes (Kovashka, Parikh, and Grauman 2012) (it does not contain any predefined split, thus we split it in a training set of 7.4K and a test set of 7.3K images). Clothing Attributes Dataset (Chen, Gallagher, and Girod 2012) (CAD) with around 2K images partially annotated for 7 clothing categories, 23 binary and 2 multi-class attributes (due to its very small size and to the non-predefined split, we perform 6 times 2-fold cross validation, i.e., we create 6 different 50-50 splits of the data).
@@
Case Study I: We used RetinaFace (Deng et al. 2020) to extract bboxes and 5 facial landmarks (for alignment) and resized images to $112\times 112\times 3$. Mixaugment (Psaroudakis and Kollias 2022) was used for data augmentation. Case Study II: for CelebA we used the database’s aligned images and resized them to $112\times 112\times 3$ (batch size = 200, Adam, lr = $10^{-3}$); for CUB, we cropped the images using the bboxes, resized them to $280\times 280\times 3$, we used label smoothing (value = 0.3) and performed affine transformations as data augmentation (batch size = 150, Adam, lr = $10^{-3}$); for S-ADD, we resized images to $280\times 280\times 3$ (batch size = 100, Adam, lr = $10^{-3}$); for CAD, we resized images to $280\times 280\times 3$ (batch size = 50, Adam, lr = $10^{-3}$). In all databases, images were normalized to $[-1,1]$. Tesla V100 32GB GPU & Tensorflow were used for model training.